Solve: (x/x+2)- (4/x-4)=(1/2)
Answer by edjones(8007)

(x/x+2)-(4/x-4)=(1/2)
2x(x-4)-8(x+2)=(x+2)(x-4) Multiply each side by 2(x-4)(x+2) to eliminate fractions.
2x^2-8x-8x-16=x^2-2x-8
x^2-14x-8=0 Subtract the left side from the right.
x^2-14x =8
x^2-14x+49=8+49 complete the square.
(x-7)^2=57
x-7=+-sqrt(57) sqrt of each side.
x=7+sqrt(57), x=7-sqrt(57)
.
Ed